Suppose that the economy is characterised by the following behavioural equations: C=C0+c1Yd where C0=280
I=270
G=300
T=200
Marginal Propensity to Save (MPS) is 0,4 or 40% solve for:
a) equlibrium GDP(Y)
b) disposal income (Yd)
c) consumption spending(C)
1
Expert's answer
2020-04-23T12:24:41-0400

"Y= C + I + G - T"

MPC = 1 - 0.4 = 0.6

C = 280 + 0.6Yd

GDP(Y) = (280 + 0.6Yd) + 270 + 300 - 200

Y = 650 + 0.6Yd